OpenStudy (anonymous):

How much pure acid should be mixed with 3 gallons of a 20% acid solution in order to get a 40% acid solution?

OpenStudy (anonymous):

100 y+(20*3)=40(y+3) 100y +60 = 40y +120 100y-40y =120-60 60y = 60 y=1 need 1 gallons

OpenStudy (amistre64):

3(.20) + b(1) = 3(.40) + b(.40) b(1-.40) = 3(.40-.20) i get 1 as well